OpenStudy (anonymous):

754-543=? I need compensation done, please!

OpenStudy (anonymous):

use a calculator. 211.

OpenStudy (anonymous):

Or you could do it in your head easily. 7-5=2 5-4=1 4-3=1 =211

OpenStudy (anonymous):

No i need compensation as if u -3 off one then you need to +3 to the other! I can't just use a calculator!

OpenStudy (anonymous):

 I hope that helped :)

OpenStudy (anonymous):

add 57 to the first number and subtract 57 from the second to get 811 - 600 = 211

OpenStudy (anonymous):

Well I need to compensate Example, if you -3 off one number then you have to +3 to the other number to make it easier to add or subtract mentally!